# Zoom-IQA: Image Quality Assessment with Reliable Region-Aware Reasoning

### v1

@@ 3.2 Self-Guided Exploration
KL-Coverage Regularizer. Recent studies on applying Reinforcement Learning (RL) to large language models (LLMs) [cui2025entropy, cheng2025reasoning] highlight a critical challenge: policy entropy often drops sharply at the onset of training, declining monotonically to near zero. This “entropy collapse” severely limits the model’s ability to explore, leading to performance plateaus. In the context of IQA, this issue is particularly detrimental. It leads to a collapse in the diversity of both reasoning paths and predicted rating scores. For instance, existing RL-based IQA methods, such as Visualquality-R1 [wu2025visualquality], suffer from “score collapse.” On the KonIQ [hosu2020koniq] test set, this method’s output unique score ratio is merely 2.04%, in stark contrast to the 71.34% of the ground-truth Mean Opinion Scores (MOS) distribution (when rounded to two decimal places).
To address this problem, we propose the KL-Coverage regularizer. This approach is inspired by the use of KL penalties to constrain policy updates [schulman2017proximal] and the recent finding that high covariance between action log-probabilities and logit changes leads to rapid policy entropy collapse [cui2025entropy]. Our regularizer is thus designed to specifically suppress numerical tokens that exhibit this high covariance
@@
Figure 4: Qualitative comparison of Zoom-IQA with competing methods (Q-insight [li2025q], VisualQuality-R1 [wu2025visualquality]). We highlight: correct descriptions, incorrect descriptions, and the uncertainty-aware reasoning unique to our model.
Category Methods KonIQ SPAQ KADID PIPAL LiveW AGIQA CSIQ
Handcrafted NIQE [mittal2012making] 0.533 / 0.530 0.679 / 0.664 0.468 / 0.405 0.195 / 0.161 0.493 / 0.449 0.560 / 0.533 0.718 / 0.628
BRISQUE [mittal2012no] 0.225 / 0.226 0.490 / 0.406 0.429 / 0.356 0.267 / 0.232 0.361 / 0.313 0.541 / 0.497 0.740 / 0.556
Non-VLM NIMA [talebi2018nima] 0.896 / 0.859 0.838 / 0.856 0.532 / 0.535 0.390 / 0.399 0.814 / 0.771 0.715 / 0.654 0.695 / 0.649
HyperIQA [su2020blindly] 0.917 / 0.906 0.791 / 0.788 0.506 / 0.468 0.410 / 0.403 0.772 / 0.749 0.702 / 0.640 0.752 / 0.717
DBCNN [zhang2018blind] 0.884 / 0.875 0.812 / 0.806 0.497 / 0.484 0.384 / 0.381 0.773 / 0.755 0.730 / 0.641 0.586 / 0.572
MUSIQ [ke2021musiq] 0.924 / 0.929 0.868 / 0.863 0.575 / 0.556 0.431 / 0.431 0.789 / 0.830 0.722 / 0.630 0.771 / 0.710
ManIQA [yang2022maniqa] 0.849 / 0.834 0.768 / 0.758 0.499 / 0.465 0.457 / 0.452 0.849 / 0.832 0.723 / 0.636 0.623 / 0.627
VLM (w/o & w/ reasoning) CLIP-IQA+ [wang2022exploring] 0.909 / 0.895 0.866 / 0.864 0.653 / 0.654 0.427 / 0.419 0.832 / 0.805 0.736 / 0.685 0.772 / 0.719
C2Score [zhu2024adaptive] 0.923 / 0.910 0.867 / 0.860 0.500 / 0.453 0.354 / 0.342 0.786 / 0.772 0.777 / 0.671 0.735 / 0.705
Q-Align [wu2023q] 0.941 / 0.940 0.886 / 0.887 0.674 / 0.684 0.403 / 0.419 0.853 / 0.860 0.772 / 0.735 0.671 / 0.737
DeQA [you2025teaching] 0.953 / 0.941 0.895 / 0.896 0.694 / 0.687 0.472 / 0.478 0.892 / 0.879 0.809 / 0.729 0.787 / 0.744
Q-Insight [li2025q] 0.918 / 0.895 0.903 / 0.903 0.702 / 0.702 0.458 / 0.435 0.870 / 0.839 0.816 / 0.766 0.685 / 0.640
VisualQuality-R1 [wu2025visualquality] 0.910 / 0.896 0.889 / 0.892 0.703 / 0.712 0.451 / 0.441 0.856 / 0.827 0.817 / 0.760 0.768 / 0.707
@@
Rank Reward. To ensure the model learns relative quality ordering, we define a rank reward $R_{\text{rank}}(x_{i})$ based on pairwise comparisons within a batch, inspired by the Thurstone model [thurstone2017law]. It is computed as:
$R_{\text{rank}}(x_{i})=\frac{1}{B-1}\sum_{j\neq i}\left(\sqrt{\hat{p}_{ij}p^{*}_{ij}}+\sqrt{(1-\hat{p}_{ij})(1-p^{*}_{ij})}\right).$ (5)
@@
DepictQA [you2024depicting] 5.40 5.49 5.51 7.96 4.54 5.09 4.41 7.80 6.04 6.39 6.14 7.86 4.80 5.08 4.46 6.87
VisualQuality-R1 [wu2025visualquality] 7.29 7.60 7.29 7.57 6.10 6.05 6.02 6.79 8.32 8.35 7.70 7.55 6.61 6.67 5.51 6.82
Q-Insight [li2025q] 7.17 7.44 7.08 6.93 5.32 5.74 5.29 6.15 7.84 8.02 7.51 6.98 6.22 6.35 5.54 5.90
@@ 4.1 Experimental Settings
Implementation Details. We initialize Qwen2.5-VL-7b [bai2025qwen2] as our base model during the first cold start stage, in which training is performed with a batch size of 2, 8 gradient accumulation steps, a learning rate of $2.5\times 10^{-6}$, and a warm-up ratio of 0.3. For GRPO, we train the finetuned model after the cold start stage with a batch size of 1, 2 gradient accumulation steps, a learning rate of $1\times 10^{-6}$, and a KL penalty coefficient of $\beta=0.04$. The number of generated responses N is set to 8.
Datasets and Metrics. For the first cold start stage, we applied SFT with our collected high-quality CoT datasets using cross-entropy loss. For the score regression task, we conduct training and evaluation on six IQA datasets grouped into three categories: (1) In-the-wild datasets, including KonIQ [hosu2020koniq], SPAQ [fang2020perceptual], and LIVE-Wild [ghadiyaram2015live]; (2) Synthetic distortion datasets, including KADID [lin2019kadid], PIPAL [jinjin2020pipal], and CSIQ [larson2010most]; (3)AI-generated image datasets, including AGIQA [li2023agiqa]. We adopt the pearson linear correlation coefficient (PLCC) and spearman rank-order correlation coefficient (SRCC) as metrics to evaluate performance on the score regression task, following previous works [you2025teaching, li2025q].
@@ 4.2 Comparison and Evaluation
Image Quality Score Regression. We compare our method with SOTA IQA methods in three different categories: (I) handcrafted, including NIQE [mittal2012making] and BRISQUE [mittal2012no] ; (II) deep learning-based, NIMA [talebi2018nima], HyperIQA [su2020blindly], DBCNN [zhang2018blind], MUSIQ [ke2021musiq], and ManIQA [yang2022maniqa] (III) MLLM-based models, CLIP-IQA+ [wang2023exploring], C2Score [zhu2024adaptive], Q-Align [wu2023q], DeQA-Score [you2025teaching], and Q-Insight [li2025q]. Since VisualQualityR1 [wu2025visualquality] did not report a KonIQ-only trained model, we retrained it with its official training code. As shown in Table 1, our approach achieves comparable performance compared with existing baselines across various synthetic and real-world benchmarks. When comparing with state-of-the-art IQA methods [li2025q, wu2025visualquality] w/ reasoning capability, our Zoom-IQA presents consistently superior performance across almost all the benchmarks. Furthermore, a qualitative comparison demonstrates the superiority of our region-aware reasoning over competing methods (Fig. 4).
@@
Image Quality Reasoning. To validate the effectiveness and accuracy of our reasoning chains, we follow common practices [chen2024mllm, jiang2025mme] to employ a VLM-as-judge evaluation methodology on the KonIQ and SPAQ datasets. We utilize two powerful, closed-source VLMs (Gemini-2.5-Flash [comanici2025gemini] and GPT-5-mini) as evaluators, which are tasked to score the generated descriptions on a 1-to-9 scale across four key criteria: Accuracy, Reasonableness, Completeness, and Confidence. To ground the assessments and ensure objectivity, the VLMs are prompted with the image, the generated reasoning chain, and corresponding low-level image indicators (e.g., brightness, sharpness) for cross-referencing. The detailed definitions of each metric and the full prompt structure are provided in the Appendix. As shown in Table 2, our method (Zoom-IQA) consistently and significantly outperforms all baselines [you2024depicting, wu2025visualquality, li2025q] across both datasets and under the scrutiny of both VLM evaluators, indicating the superiority of our reasoning reliability.
Reasoning-guided Restoration. High-quality reasoning should be able to provide reliable guidance for downstream tasks such as restoration [wang2024exploiting, yu2024scaling, wang2025seedvr]. To further demonstrate the superiority of Zoom-IQA’s reasoning capability, we extract IQA reasoning as guidance for SUPIR [yu2024scaling], a state-of-the-art restoration model that accepts textual prompts. We replace its default prompt generator (LLaVA-1.5-13b [liu2023visual]) and instead feed it the textual reasoning outputs from Q-Insight [li2025q], VisualQuality-R1 [wu2025visualquality], and our own Zoom-IQA to examine how IQA reasoning affects restoration outputs. As visualized in Figure 5, the reasoning generated by our method provides more appropriate and specific guidance, enabling SUPIR to restore fine-grained, detailed textures from the input images, which are overlooked or poorly reconstructed when guided by the reasoning from other methods. Such a comparison demonstrates the practical effectiveness of our model’s reasoning, highlighting the reliability of our IQA reasoning.
@@ 6.1 Data Construction Prompts
Tab. 4 and 5 detail the prompts used to construct our Grounded-Rationale-IQA (GR-IQA) dataset with Gemini-2.5-pro [comanici2025gemini]. The thinking rationale comprises four components:
@@ 6.2 Visual Reliance Filtering
Due to the limited API access of closed-source VLMs, we cannot compute the log-probabilities for predefined text sequences (or specific candidate answers). Consequently, existing hallucination detection methods relying on offline contrastive probability analysis [xing2025scalecap] are inapplicable. Furthermore, prior online decoding strategies [leng2024mitigating, huo2024self] typically inject image distortions to verify consistency. While effective for semantic-level tasks—where content identity remains robust to noise—this approach is fundamentally incompatible with IQA. Since IQA aims to precisely evaluate visual degradation, introducing artificial distortion alters the target attribute itself. This makes it difficult to disentangle whether the model is responding to the original image artifacts or the injected noise.
@@ 6.3 Hint-Augmented Consistency Filtering
Tab. 6 presents examples of data discarded by our Hint-Augmented Consistency Filtering with Qwen-2.5-32b [qwen2.5]. Specifically, we leverage the low-level hints from KonIQ [hosu2020koniq] and the quality hints from KonIQ++ [su2021koniq++] as reference; this allows our method to effectively filter out generated prompts that are inconsistent with human labels.
@@ 7 Image Quality Reasoning
Tab. 7 presents the prompts used to validate the effectiveness of our reasoning chains on the KonIQ and SPAQ [fang2020perceptual] datasets. These datasets provide low-level image attributes (brightness, contrast, colorfulness, and sharpness for KonIQ; brightness, colorfulness, contrast, noisiness, and sharpness for SPAQ) alongside MOS scores, which ensures more precise evaluation judgments.
@@ 8 Reasoning-guided Restoration
To further demonstrate the efficacy of Zoom-IQA’s reasoning capability, we evaluate its performance on the task of text-guided image restoration. We employ the DreamClear [ai2024dreamclear] framework for this experiment, primarily because its T5 text encoder [raffel2020exploring] supports a significantly longer context window than the CLIP encoder [radford2021learning] used in SUPIR [yu2024scaling], thereby accommodating the detailed reasoning prompts generated by our method. Specifically, we adhere to the standard DreamClear pipeline, where initial restoration results from a lightweight network (SwinIR [liang2021swinir]) are input into VLMs to generate guidance. We conduct experiments under three distinct prompt settings. In addition to the standard captions from LLaVA-1.6-13b [liu2023visual] used by the original DreamClear, we separately employ the reasoning content from VisualQuality-R1 and Q-Insight, and the “Directions for Improvement” derived from our method’s thinking process. As illustrated in Figs. 7-9, restoration guided by our method’s reasoning exhibits superior perceptual texture quality, particularly in complex regions such as facial features (Fig. 8) and fur (Fig. 9).
@@ 9 More Qualitative Comparison Results
We provide additional qualitative comparisons between our method (Zoom-IQA) and competing methods (Q-Insight [li2025q], VisualQuality-R1 [wu2025visualquality]) on real-world images, as illustrated in Fig. 10 through Fig. 13. These results demonstrate the superiority of our reasoning mechanism, which not only identifies specific distortions but also explicitly localizes the distorted objects/regions. Furthermore, in complex scenes (e.g., Fig. 10 and Fig. 11), our method employs interactive, region-aware reasoning: it first hypothesizes potential flaws (green text), then grounds them via adaptive cropping (orange text), and finally verifies the degradation (blue text). This hypothesize-and-verify loop ensures a comprehensive assessment. Conversely, in scenes with simpler compositions (e.g., Fig. 12 and Fig. 13), our method directly detects global distortions without performing cropping operations.

### v2

@@ 3.2 Self-Guided Exploration
KL-Coverage Regularizer. Recent studies on applying Reinforcement Learning (RL) to large language models (LLMs) [13, 10] highlight a critical challenge: policy entropy often drops sharply at the onset of training, declining monotonically to near zero. This “entropy collapse” severely limits the model’s ability to explore, leading to performance plateaus. In the context of IQA, this issue is particularly detrimental. It leads to a collapse in the diversity of both reasoning paths and predicted rating scores. For instance, existing RL-based IQA methods, such as Visualquality-R1 [67], suffer from “score collapse.” On the KonIQ [18] test set, this method’s output unique score ratio is merely 2.04%, in stark contrast to the 71.34% of the ground-truth Mean Opinion Scores (MOS) distribution (when rounded to two decimal places). To address this problem, we propose the KL-Coverage regularizer. This approach is inspired by the use of KL penalties to constrain policy updates [47] and the recent finding that high covariance between action log-probabilities and logit changes leads to rapid policy entropy collapse [13]. Our regularizer is thus designed to specifically suppress numerical tokens that exhibit this high covariance.
@@
Figure 4: Qualitative comparison of Zoom-IQA with competing methods (Q-insight [29], VisualQuality-R1 [67]). We highlight: correct descriptions, incorrect descriptions, and the uncertainty-aware reasoning unique to our model.
@@
Rank Reward. To ensure the model learns relative quality ordering, we define a rank reward $R_{\text{rank}}(x_{i})$ based on pairwise comparisons within a batch, inspired by the Thurstone model [57]. It is computed as:
$R_{\text{rank}}(x_{i})=\frac{1}{B-1}\sum_{j\neq i}\left(\sqrt{\hat{p}_{ij}p^{*}_{ij}}+\sqrt{(1-\hat{p}_{ij})(1-p^{*}_{ij})}\right),$ (5)
@@
Category Methods KonIQ SPAQ KADID PIPAL LiveW AGIQA CSIQ
Handcrafted NIQE [40] 0.533 / 0.530 0.679 / 0.664 0.468 / 0.405 0.195 / 0.161 0.493 / 0.449 0.560 / 0.533 0.718 / 0.628
BRISQUE [39] 0.225 / 0.226 0.490 / 0.406 0.429 / 0.356 0.267 / 0.232 0.361 / 0.313 0.541 / 0.497 0.740 / 0.556
Non-VLM NIMA [56] 0.896 / 0.859 0.838 / 0.856 0.532 / 0.535 0.390 / 0.399 0.814 / 0.771 0.715 / 0.654 0.695 / 0.649
HyperIQA [53] 0.917 / 0.906 0.791 / 0.788 0.506 / 0.468 0.410 / 0.403 0.772 / 0.749 0.702 / 0.640 0.752 / 0.717
DBCNN [77] 0.884 / 0.875 0.812 / 0.806 0.497 / 0.484 0.384 / 0.381 0.773 / 0.755 0.730 / 0.641 0.586 / 0.572
MUSIQ [25] 0.924 / 0.929 0.868 / 0.863 0.575 / 0.556 0.431 / 0.431 0.789 / 0.830 0.722 / 0.630 0.771 / 0.710
ManIQA [70] 0.849 / 0.834 0.768 / 0.758 0.499 / 0.465 0.457 / 0.452 0.849 / 0.832 0.723 / 0.636 0.623 / 0.627
VLM (w/o & w/ reasoning) CLIP-IQA+ [59] 0.909 / 0.895 0.866 / 0.864 0.653 / 0.654 0.427 / 0.419 0.832 / 0.805 0.736 / 0.685 0.772 / 0.719
C2Score [81] 0.923 / 0.910 0.867 / 0.860 0.500 / 0.453 0.354 / 0.342 0.786 / 0.772 0.777 / 0.671 0.735 / 0.705
Q-Align [65] 0.941 / 0.940 0.886 / 0.887 0.674 / 0.684 0.403 / 0.419 0.853 / 0.860 0.772 / 0.735 0.671 / 0.737
DeQA [71] 0.953 / 0.941 0.895 / 0.896 0.694 / 0.687 0.472 / 0.478 0.892 / 0.879 0.809 / 0.729 0.787 / 0.744
Q-Insight [29] 0.918 / 0.895 0.903 / 0.903 0.702 / 0.702 0.458 / 0.435 0.870 / 0.839 0.816 / 0.766 0.685 / 0.640
VisualQuality-R1 [67] 0.910 / 0.896 0.889 / 0.892 0.703 / 0.712 0.451 / 0.441 0.856 / 0.827 0.817 / 0.760 0.768 / 0.707
@@
DepictQA [73] 5.40 5.49 5.51 7.96 4.54 5.09 4.41 7.80 6.04 6.39 6.14 7.86 4.80 5.08 4.46 6.87
VisualQuality-R1 [67] 7.29 7.60 7.29 7.57 6.10 6.05 6.02 6.79 8.32 8.35 7.70 7.55 6.61 6.67 5.51 6.82
Q-Insight [29] 7.17 7.44 7.08 6.93 5.32 5.74 5.29 6.15 7.84 8.02 7.51 6.98 6.22 6.35 5.54 5.90
@@ 4.1 Experimental Settings
Implementation Details. We initialize Qwen2.5-VL-7b [2] as our base model during the first cold start stage, in which training is performed with a batch size of 2, 8 gradient accumulation steps, a learning rate of $2.5\times 10^{-6}$, and a warm-up ratio of 0.3. For GRPO, we train the finetuned model after the cold start stage with a batch size of 1, 2 gradient accumulation steps, a learning rate of $1\times 10^{-6}$, and a KL penalty coefficient of $\beta=0.04$. The number of generated responses N is set to 8.
Datasets and Metrics. For the first cold start stage, we applied SFT with our collected high-quality CoT datasets using cross-entropy loss. For the score regression task, we conduct training and evaluation on six IQA datasets grouped into three categories: (1) In-the-wild datasets, including KonIQ [18], SPAQ [15], and LIVE-Wild [16]; (2) Synthetic distortion datasets, including KADID [31], PIPAL [23], and CSIQ [26]; (3)AI-generated datasets, including AGIQA [28]. We adopt the pearson linear correlation coefficient (PLCC) and spearman rank-order correlation coefficient (SRCC) as metrics to evaluate performance on the score regression task, following previous works [71, 29].
@@ 4.2 Comparison and Evaluation
Image Quality Score Regression. We compare our method with SOTA IQA methods in three different categories: (I) handcrafted, including NIQE [40] and BRISQUE [39] ; (II) deep learning-based, NIMA [56], HyperIQA [53], DBCNN [77], MUSIQ [25], and ManIQA [70] (III) MLLM-based models, CLIP-IQA+ [60], C2Score [81], Q-Align [65], DeQA-Score [71], and Q-Insight [29]. Since VisualQualityR1 [67] did not report a KonIQ-only trained model, we retrained it with its official training code. As shown in Table 1, our approach achieves comparable performance compared with existing baselines across various synthetic and real-world benchmarks. When comparing with state-of-the-art IQA methods [29, 67] w/ reasoning capability, our Zoom-IQA presents consistently superior performance across almost all the benchmarks. Furthermore, a qualitative comparison demonstrates the superiority of our region-aware reasoning over competing methods (Fig. 4).
@@
Image Quality Reasoning. To validate the effectiveness and accuracy of our reasoning chains, we follow common practices [7, 22] to employ a VLM-as-judge evaluation methodology on the KonIQ and SPAQ datasets. We utilize two powerful, closed-source VLMs (Gemini-2.5-Flash [12] and GPT-5-mini) as evaluators, which are tasked to score the generated descriptions on a 1-to-9 scale across four key criteria: Accuracy, Reasonableness, Completeness, and Confidence. To ground the assessments and ensure objectivity, the VLMs are prompted with the image, the generated reasoning chain, and corresponding low-level image indicators (e.g., brightness, sharpness) for cross-referencing. The detailed definitions of each metric and the full prompt structure are provided in the Appendix. As shown in Table 2, our method (Zoom-IQA) consistently and significantly outperforms all baselines [73, 67, 29] across both datasets and under the scrutiny of both VLM evaluators, indicating the superiority of our reasoning reliability.
Reasoning-guided Restoration. High-quality reasoning should be able to provide reliable guidance for downstream tasks such as restoration [62, 75, 61]. To further demonstrate the superiority of Zoom-IQA’s reasoning capability, we extract IQA reasoning as guidance for SUPIR [75], a state-of-the-art restoration model that accepts textual prompts. We replace its default prompt generator (LLaVA-1.5-13b [32]) and instead feed it the textual reasoning outputs from Q-Insight [29], VisualQuality-R1 [67], and our own Zoom-IQA to examine how IQA reasoning affects restoration outputs. As visualized in Figure 5, the reasoning generated by our method provides more appropriate and specific guidance, enabling SUPIR to restore fine-grained, detailed textures from the input images, which are overlooked or poorly reconstructed when guided by the reasoning from other methods. Such a comparison demonstrates the practical effectiveness of our model’s reasoning, highlighting the reliability of our IQA reasoning.
@@ 6.1 Data Construction Prompts
Tab. 4 and 5 detail the prompts used to construct our Grounded-Rationale-IQA (GR-IQA) dataset with Gemini-2.5-pro [12]. The thinking rationale comprises four components:
@@ 6.2 Visual Reliance Filtering
Due to the limited API access of closed-source VLMs, we cannot compute the log-probabilities for predefined text sequences (or specific candidate answers). Consequently, existing hallucination detection methods relying on offline contrastive probability analysis [68] are inapplicable. Furthermore, prior online decoding strategies [27, 20] typically inject image distortions to verify consistency. While effective for semantic-level tasks—where content identity remains robust to noise—this approach is fundamentally incompatible with IQA. Since IQA aims to precisely evaluate visual degradation, introducing artificial distortion alters the target attribute itself. This makes it difficult to disentangle whether the model is responding to the original image artifacts or the injected noise.
@@ 6.3 Hint-Augmented Consistency Filtering
Tab. 6 presents examples of data discarded by our Hint-Augmented Consistency Filtering with Qwen-2.5-32b [69]. Specifically, we leverage the low-level hints from KonIQ [18] and the quality hints from KonIQ++ [52] as reference; this allows our method to effectively filter out generated prompts that are inconsistent with human labels.
@@ 7 Image Quality Reasoning
Tab. 7 presents the prompts used to validate the effectiveness of our reasoning chains on the KonIQ and SPAQ [15] datasets. These datasets provide low-level image attributes (brightness, contrast, colorfulness, and sharpness for KonIQ; brightness, colorfulness, contrast, noisiness, and sharpness for SPAQ) alongside MOS scores, which ensures more precise evaluation judgments.
@@ 8 Reasoning-guided Restoration
To further demonstrate the efficacy of Zoom-IQA’s reasoning capability, we evaluate its performance on the task of text-guided image restoration. We employ the DreamClear [1] framework for this experiment, primarily because its T5 text encoder [45] supports a significantly longer context window than the CLIP encoder [44] used in SUPIR [75], thereby accommodating the detailed reasoning prompts generated by our method. Specifically, we adhere to the standard DreamClear pipeline, where initial restoration results from a lightweight network (SwinIR [30]) are input into VLMs to generate guidance. We conduct experiments under three distinct prompt settings. In addition to the standard captions from LLaVA-1.6-13b [32] used by the original DreamClear, we separately employ the reasoning content from VisualQuality-R1 and Q-Insight, and the “Directions for Improvement” derived from our method’s thinking process. As illustrated in Figs. 7-9, restoration guided by our method’s reasoning exhibits superior perceptual texture quality, particularly in complex regions such as facial features (Fig. 8) and fur (Fig. 9).
@@ 9 More Qualitative Comparison Results
We provide additional qualitative comparisons between our method (Zoom-IQA) and competing methods (Q-Insight [29], VisualQuality-R1 [67]) on real-world images, as illustrated in Fig. 10 through Fig. 13. These results demonstrate the superiority of our reasoning mechanism, which not only identifies specific distortions but also explicitly localizes the distorted objects/regions. Furthermore, in complex scenes (e.g., Fig. 10 and Fig. 11), our method employs interactive, region-aware reasoning: it first hypothesizes potential flaws (green text), then grounds them via adaptive cropping (orange text), and finally verifies the degradation (blue text). This hypothesize-and-verify loop ensures a comprehensive assessment. Conversely, in scenes with simpler compositions (e.g., Fig. 12 and Fig. 13), our method directly detects global distortions without performing cropping operations.